please explain step by step, that is the right answer
X 8.5.1 Find the partial fraction decomposition for the rational expression. 9x + 38 (x + 6)(x + 2)
Answer #1

9x+38 x+6 +2 (x+6) (2+2) A= - 9(-6) +38 -6+2 9x+38 - x+2 - 4 /a=-1 9x+38 x+6 la=-2 9(-2) +385 -2+6 at the OCT6 T x+2.